OpenStudy (anonymous):

A tortoise is walking in the desert. It walks at a speed of 3 meters per minute for 37.5 meters. For how many minutes does it walk?

OpenStudy (anonymous):

helppp

OpenStudy (anonymous):

Speed=distance divided by time =d/t 3m/min=37.5m/t t=(37.5m)/(3m/min) the meters goes away and you left with t in min t=12.5 min

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

what about this one? 2x+17y=19